A 30-kg solid sphere of radius 0.12 m is rolling without slipping on a horizontal surface at 1.7 m/s. What average torque is required to stop the sphere in 5.0 rev without inducing skidding?
Answer #1

I =( 2/5) MR2 = 0.4 *30 *0.0144 = 0.173

Velocity = 1.7 m/s.

W = 1.7/0.12 = 14.17 rad/sec

Angle of rotation = 5*2pi = 10*3.14 = 31.4 radian

0 = 14.172 - 2*a*31.4

a = 3.2 rad/s2

Now Average torque = Ia = 0.173 * 3.2 = 5.53 Nm